OpenStudy (anonymous):

Two objects whose masses are 6.0kg and 11.0kg are suspended by strings. Find the tension.

OpenStudy (ash2326):

T=mg T1=6*10=60 Newtons T2=11.0*10=110 Newtons

OpenStudy (hunus):

 Because the two objects are stationary it can be said that they have a net zero force. Because the only force acting on them is the force of gravity, the tension in the strings must be equal to the force of gravity on each object. The force of gravity is given by the equation \[f_{g} = m*a\] where a is gravity. so the tension in the 6kg object is given by \[(6kg)(10 m/s^2)=60N\] And the tension in the 11kg object is given by \[(11kg)(10m/s^2)=110N\]